Make New Entry,
Make Followup Entry
User name Markowitz
Log entry time 08:08:35 on July 07, 2010
Entry number 330614
keyword=Owl shift summary
Owl Shift Summary: July 7, 2010
- Pete Markowitz
- Victor Maxwell
- Eric Jensen
- 00:00 Sieve is in, T6 trigger is modified, taking data on Ta
target. Run 1815 underway, left arm data for optics.
- 00:20 Ask for access to put sieve half-out. Turn off septum, put
both sieves half-out, exit, turn on septum, ask for 20 uA
- 00:45 Start run 1816, raster off, Ta target, modified T6 triggers
(only RHRS S2m && RHRS GC)
- 00:58 Ask for access to take sieve out. Turn off septum, put both
sieves out, exit, turn on septum, ask to start at 10 uA.
- 01:40 MCC calls. South arc tripped off and the magnets will have
to be cycled. Beam will be gone for ~20 minutes.
- 02:13 Start run 1818, raster off, sieve out. PID, ~1KHz T3, 250
Hz T6 modified.
- 02:33 Start run 1819, raster off, sieve out, T1.
- 02:40 Ask for access to restore T6 trigger. Septum off.
- Runs 1820, 1821, and 1822 are junk runs. We adjusted the raster
to avoid the edge of anything and check rates for prescales.
- 03:36 Start run 1823, production. Raster 8x6.
- 04:19 Start run 1824, production. Raster 8x6.
- 04:53 Start run 1825, production. Raster 8x6.
- 05:42 Start run 1826, production. Raster 8x6.
- 06:52 Start run 1827, production. Raster 8x6.
- 08:03 Start run 1828, production. Raster 8x6.